Classify and reduce the partial differential equation to it's canonical form

Uxx+U2xy+Uyy =0


1
Expert's answer
2021-05-07T10:00:58-0400

Given PDE is-

"U_{xx}+2U_{xy}+U_{yy}=0"


or, "\dfrac{d^2U}{dx^2}+\dfrac{2d^2U}{dxdy}+\dfrac{d^2U}{dy^2}=0"


i.e ". r+2s+t=0"


Then The required canonical form is-

"rR+sS+tT+\lambda(x,y,z,p,q)=0"


Where "R=1,S=2,T=1"


"S^2-4RT=(2)^2-4(1)(1)=0"

Given Equation is parabolic.


Also, "R\lambda^2+S\lambda+T=0"

"\lambda^2+2\lambda+1=0\\(\lambda+1)^2=0\\\lambda=-1,-1"


Then the characterstics equation is-

"\dfrac{dy}{dx}+\lambda=0\\

 \dfrac{dy}{dx}-1=0\\dy-dx=0"


Integrating both the sides and we get the required canonical form as-

"y-x=c"